Hennops Ltd has made a profit after tax of R23.4 million in their current financial year, with a commensurate growth in their cash reserves of R26.1 million. They are considering a new capital project of R28 million, which the operations director has suggested be funded as follows: R14 million in cash and R14 million by way of a long-term loan with a fixed interest rate of 15% per annum.

The financial director is aware that Hennops Ltd should continue paying a dividend, as not doing so may impact negatively on their current share price of 800 cents per share, trading at twice the current net asset value. Hennops Ltd paid a dividend of 30 cents per share in respect of their previous financial year. Their 15 million shares in issue (in terms of number) have remained unchanged for the last three years.

Required:

The financial director is considering a number of dividend options for the current year:

1.1 Growing the previous dividend by 10%.

1.2 Applying a dividend cover of 4.

1.3 Issuing a script dividend on a 1 for 20 basis assuming all shareholders accepts the offer.

1.4 Repurchase 1 million shares at the current market price on a pro-rata basis.

Show the impact of each of the above per share and on Hennops Ltd’s statement of Financial Position
Answer #1
1.1 Growing dividend by 10%
Dividend =30*1.1=33 cents
Total dividend=15million*33 cents            4,950,000 RM
JOURNAL ENTRY
ACCOUNT DEBIT CREDIT
Retained Earnings            4,950,000
Cash    4,950,000
CASH RESERVE WILL DECREASE BY 4.95 Million
Cash reserve=(26.1 -4.95 )million          21,150,000 RM
Dividend per share=(4.95/15)                       0.33 RM
Cash Reserve per share=21150000/15000000                       1.41 RM
1.2 Dividend Cover of 4
Profit after tax =23.4 million
Total dividend=(23.4 /4)million            5,850,000 RM
JOURNAL ENTRY
ACCOUNT DEBIT CREDIT
Retained Earnings            5,850,000
Cash    5,850,000
CASH RESERVE WILL DECREASE BY 5.85 Million
Cash reserve=(26.1 -5.85 )million          20,350,000 RM
Dividend per share=(5.85/15)                       0.39 RM
Cash Reserve per share=20350000/15000000                       1.36 RM
1.3 SCRIPT DIVIDEND ON 1 for 20 Basis
Number of share outstanding=15 million
Number of shares in script dividend =(15/20)million                750,000
Value of Dividend   =750000*800 cents            6,000,000
JOURNAL ENTRY
ACCOUNT DEBIT CREDIT
Retained Earnings            6,000,000
Stockholders Equity    6,000,000
No Cash will be paid
There will be no change in Cash Reserve
Cash Reserve          26,100,000 RM
Cash Reserve per share=26100000/(15000000+750000) 1.66 RM
1.4 REPURCHASE 1 million shares
Total payment for repurchase =1million*800 cents            8,000,000 RM
JOURNAL ENTRY
ACCOUNT DEBIT CREDIT
Treasury Srock            8,000,000
Cash    8,000,000
Cash Reserve will decrease by 8 million
Cash Reserve          18,100,000 RM
Cash Reserve per share=18100000/(15000000-1000000)                       1.29 RM